OpenStudy (anonymous):

Stable High School has a total of 112 boys and girls who play sports. If the number of boys, b is 16 more than twice the number of girls, g, how many boys play sports at this high school? Which system of equations would help you solve this problem? 1. b + g = 112 b = 2g + 16 2. b + g = 112 g = 2b + 16 3. b − g = 112 b = 2g − 16 4. None of the systems above would solve this problem.

OpenStudy (anonymous):

Number of boys = b Number of girls = g Number of girls is 16 more than twice number of boys g=2b+16 112 boys and girls b+g=112 Substitute the g: b+2b+16=112 3b+16=112 3b=112-16 3b=94 b=96/3 b=32 Number of boys is 32 Number of girls is 112-32 = 80 Hope that makes good enough sense for you. :D

OpenStudy (binary3i):

first g = 32 b= 80

OpenStudy (anonymous):

u guys r telling to different things idk which is rite LOL

OpenStudy (anonymous):

@binary3i is wrong girls = 80 boys = 32

OpenStudy (anonymous):

what number is it

OpenStudy (binary3i):

you are wrong @ZhangYan

OpenStudy (binary3i):

@ZhangYan is no of boys 16 greater than two times the no of girls. you have written wrong

OpenStudy (anonymous):

binary3i please i don't want to argue with you so....

OpenStudy (binary3i):

sorry

OpenStudy (anonymous):

@TuringTest can you check who is correct?

OpenStudy (turingtest):

". If the number of boys, b is 16 more than twice the number of girls, g" so binary is actually right you reversed the statemnent

OpenStudy (turingtest):

the number of boys, b is 16 more than twice the number of girls, g b=2g+16

OpenStudy (turingtest):

just a slight misreading of the problem; happens all the time I'm off to class, later!

OpenStudy (binary3i):

thank you @TuringTest

OpenStudy (anonymous):

haha sorry about that i must unread correctly Btw i will give you a medal for my mistake :D

OpenStudy (turingtest):

@ZhangYan you're explanation is sound though, so kudos to that :)

OpenStudy (anonymous):

:D